OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

@1800_RATCHETHUNTERS

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@Babyjoker001

OpenStudy (anonymous):

@Catseyeglint911

OpenStudy (anonymous):

f(3) = f(2) + f(1) => 3 + 2 = 5 f(4) = f(3) + f(2) => 5 + 3 = 8 f(5) = f(4) + f(3) => 8 + 5 = 13 Therefore the sequence is 2, 3, 5, 8, 13,.....a Fibonacci sequence.

OpenStudy (anonymous):

thanks again xD!!!

OpenStudy (anonymous):

no problem

OpenStudy (zarkon):

the way you wrote it f(n) = f(1) + f(2) + f(n - 1) f(3)= f(1) + f(2) + f(3 - 1)=f(1)+f(2)+f(2)=2+3+3=8

OpenStudy (zarkon):

f(4)=f(1)+f(2)+f(4-1)=f(1)+f(2)+f(3)=2+3+8=13

OpenStudy (zarkon):

f(5)=f(1)+f(2)+f(5-1)=f(1)+f(2)+f(4)=2+3+13=18